### v2

# Magnetic field-induced momentum-dependent symmetry breaking in a kagome superconductor

Abstract
When multiple degrees of freedom share similar energy scales in quantum materials, intertwined electronic orders, which exhibit broken symmetries, are often strongly coupled. Recent studies on kagome superconductors such as CsV3Sb5 report rotational and time-reversal symmetry breaking linked to a charge density wave. Here, we observe a momentum-selective response of the electronic structure of CsV3Sb5 to an external magnetic field. By performing angle-resolved photoemission spectroscopy in a tuneable magnetic field, we demonstrate that the response of the electronic structure is compatible with piezomagnetism along with strong orbital selectivity. Our results show that the origin of the time-reversal symmetry breaking is associated with the vanadium Van Hove singularities at the onset of the charge density wave order. We also demonstrate the presence of fluctuations beyond the charge ordering temperature. Our results reveal that magnetic fields can be used as tuning knobs for disentangling intertwined orders in the momentum space for quantum materials.
I Main
The kagome lattice, a prototypical two-dimensional geometrically frustrated system, has been extensively studied as a promising candidate for realizing the quantum spin liquid state Syozi (1951); Balents (2010); Han et al. (2012). Beyond spin interactions, the kagome lattice also exhibits rich physics associated with the quantum interference of the electronic hopping paths, hosting flat bands, Dirac cones, and Van Hove singularities (VHS) in its electronic structure Mielke (1991); Tanaka and Ueda (2003); Yu and Li (2012); Kiesel and Thomale (2012); Kiesel et al. (2013). When the flat bands or the VHSs are positioned near the Fermi level, electronic instabilities emerge even for relatively weak electron-electron interactions, potentially leading to the emergence of different symmetry-broken phases, such as magnetism, charge density wave (CDW), superconductivity, and electronic nematicity Norman (2016); Ko et al. (2009); Yu and Li (2012); Kiesel and Thomale (2012); Kiesel et al. (2013); Wang et al. (2013). Such a proximity to competing ordered states makes the kagome systems highly susceptible to external perturbations such as strain and magnetic field, which can be used to tune the competition among the exotic phases Liu (2020); Wang et al. (2023); Lima et al. (2023); Dey et al. (2011). Consequently, kagome metals provide an excellent platform for investigating the various intriguing symmetry-breaking phases due to the intricate interweaving among electron correlations, topology, and geometric frustration Balents (2010); Ko et al. (2009); Kang et al. (2020).
CsV3Sb5 has been identified as a kagome superconductor (Tc 3.2 K) with the V atoms forming the kagome lattice Ortiz et al. (2019, 2020). It undergoes a 222 CDW transition at 94 K, which is well captured by first-principles calculations, indicating the relevant role of the electron-phonon coupling Tan et al. (2021); Christensen et al. (2021). Interestingly, the CDW phase was also found to display unconventional properties suggestive of time-reversal symmetry-breaking (TRSB), which has been interpreted in terms of an interaction-driven loop-current phase Park et al. (2021); Lin and Nandkishore (2021); Denner et al. (2021); Christensen et al. (2022); Wagner et al. (2023); Tazai et al. (2024). These properties include an enhanced relaxation rate measured by muon spin rotation Mielke et al. (2022); Khasanov et al. (2022), magnetic-field switching of the relative intensity of the CDW Bragg peaks seen in scanning tunneling microscopy Jiang et al. (2021); Xing et al. (2024), and anisotropic and non-reciprocal electrical magneto-transport Guo et al. (2022); Wei et al. (2024); Guo et al. (2024b). On the other hand, the presence of a non-zero magnetization remains under debate, in view of contradictory reports on the existence of a spontaneous Kerr effect Xu et al. (2022); Farhang et al. (2023); Saykin et al. (2023); Wang et al. (2024). Additionally, sixfold (C6) to twofold (C2) rotational symmetry breaking (RSB) was detected at around 35 K in the CDW state of CsV3Sb5 by a combination of nuclear magnetic resonance and elastoresistance measurements Nie et al. (2022). Such a RSB was attributed to the interlayer stacking of the 222 CDW order Christensen et al. (2021) and interpreted as an electronic nematic orderXiang et al. (2021); Jin et al. (2024). While the interpretation of RSB in terms of nematic order has been challenged by recent studies that did not find a divergent nematic susceptibility Liu et al. (2024), various experimental probes in CsV3Sb5, including polarization-resolved Raman spectroscopy Wulferding et al. (2022), scanning tunneling microscopy Zhao et al. (2021); Chen et al. (2021); Li et al. (2022), and optical spectroscopy Xu et al. (2022), have consistently found RSB within the CDW state. Interestingly, resistivity anisotropy Guo et al. (2024b) and magnetic torque measurements Asaba et al. (2024) in CsV3Sb5 have proposed a connection between RSB and TRSB. Thus, to advance the understanding of this material and shed light on these conflicting experimental reports, a direct momentum-resolved spectroscopy measurement in the presence of a magnetic field is desirable.
In light of this, we utilize our recently developed magneto-ARPES technique Huang et al. (2023) to directly measure the response of the momentum-resolved electronic structure of CsV3Sb5 to an out-of-plane tunable magnetic field. ARPES has been remarkably successful in directly probing the momentum-resolved electronic spectra of quantum materials, providing a way to disentangle intertwined order parameters that exhibit distinct momentum-dependent fingerprints Damascelli et al. (2003); Sobota et al. (2021). For the AV3Sb5 (A = K, Rb, Cs) compounds, ARPES studies have revealed the key electronic characteristics inherent to the kagome lattice, along with the band folding and CDW gap associated with the CDW transition Liu et al. (2021); Luo et al. (2022); Kang et al. (2022); Hu et al. (2022). However, to date, no momentum-resolved electronic response to an external magnetic field has been reported using ARPES.
Magnetic field is commonly used as a tuning knob for other types of experiments, but it has been carefully avoided in ARPES measurements due to its unpredictable effects on the photoelectron trajectory after photoemission Ryu et al. (2023); Huang et al. (2023). Recently, we have developed a simple method for implementing an in situ tunable out-of-plane magnetic field as the sample environment during ARPES measurements via a solenoid coil (Fig. 1e and f), enabling magneto-ARPES–a term coined by an independent work developing an in-plane field for ARPES Ryu et al. (2023). While the magnetic fields available in our setup are small, below 10 mT, they are larger than the internal magnetic field values reported by SR in AV3Sb5 Mielke et al. (2022); Khasanov et al. (2022). Utilizing this capability, we performed magneto-ARPES measurements on CsV3Sb5 to probe a momentum-dependent electronic response to a magnetic field. Specifically, we observed spectral modifications to the dispersions and Fermi surfaces associated with the vanadium -orbital VHS bands around the K/K’ points of the BZ. Importantly, these spectral modifications break the C6 symmetry of the dispersion and are odd in the magnetic field, consistent with piezomagnetism Xing et al. (2024). Additionally, we observe an elliptical elongation of the antimony -orbital electron pocket around the point of the BZ in the presence of a magnetic field. In contrast to the field-induced anisotropic spectral changes observed in the vanadium -bands, which disappear above , the spectral changes around the point persists above , suggesting significant fluctuations. Our magneto-ARPES results unambiguously reveal time-reversal symmetry breaking in CsV3Sb5 interwoven with rotational symmetry breaking, and a qualitatively different magnetic-field response of the low-energy electronic states arising from the Sb -orbitals and the V -orbitals. These results not only provide strong constraints to theoretical delineation of the exotic CDW order parameter in CsV3Sb5, but also establish a new capability for disentangling momentum-dependent order parameters in the presence of a field.
II Field response of the V 3 VHS bands at K/K’
We first introduce the zero-field electronic structure of CsV3Sb5 measured using a helium lamp-based ARPES system (Fig. 1b and c). Consistent with previous reports, the Fermi surface of CsV3Sb5 is composed of a circular electron pocket around the BZ center consisting mainly of the 5 orbitals of Sb and three triangles around the BZ corner (K/K’) consisting of the 3 orbitals of V and related to the nearby VHS at the M points Kang et al. (2022); Hu et al. (2022). The measured band dispersions match well with those obtained from density-functional theory (DFT) calculations (Fig. 1c). Raw magneto-ARPES data generally consist of intrinsic effects occurring inside the sample in response to the magnetic field and extrinsic effects occurring when the field acts on the photoemitted electrons on their way to the analyzer after they are ejected from the sample. The extrinsic effects consist of a constant energy contour rotation due to the Lorentz force, photoelectron emission angle contraction, and momentum broadening (Fig. 1d-f) Huang et al. (2023). Specifically, for the small magnetic field adopted in this work (1.6 mT), only a constant energy contour rotation was discernible while the other two effects were negligible (Fig. 1g and h) Huang et al. (2023). This in-plane constant energy contour rotation is effectively energy independent within the energy range examined in this work (up to 1 eV below the Fermi level, see Supplementary Materials), allowing straightforward post-measurement correction without the need for intricate data processing. It can also be used to directly determine the field strength at the sample. As an example, Fig. 1g-i compares the Fermi surface mapped by rotating the sample measured at 0, -1.6 mT and +1.6 mT. The maps display opposite rotation directions for positive and negative fields, which can be corrected by an in-plane azimuthal angle offset before any further data analysis is performed (Fig. 1h,i). To avoid any potential complications in field-dependent comparisons arising from the photoemission matrix-element effects Damascelli et al. (2003), all subsequent measurements were performed with the DA30 deflector mode without the need to rotate or move samples when changing the magnetic field. Hence, in all comparisons that we provide with and without a field, the beam spot was kept on a fixed spot of the sample during the whole measurement, guaranteeing that the light polarization relative to the crystal axes remained the same when the field strength was changed.
We next compare the Fermi surface of CsV3Sb5 measured with and without a magnetic field in the CDW phase. To better observe the evolution of the electronic spectra in a magnetic field, we analyze the constant energy contours at -0.35 eV where the three sets of bands forming the triangular Fermi surfaces at the Fermi level () around K are well separated (Fig. 2a). In the absence of a magnetic field, two large triangular pockets ( and ) and a small circular pocket at K (K’) are resolved. The corresponding ARPES spectra are not distinguishable between K and K’ (Fig. 2b). Notably, the and sheets meet at the M point, forming two sets of "X"-shaped constant energy contours that straddle the M point across the two neighboring BZs. When a magnetic field of 1.6 mT is applied, the overall shape of the constant energy contour remains unchanged. However, one diagonal branch of the "X"-shaped sheets weakens into the spectral background while the other branch remains sharp and discernible (yellow arrow in Fig. 2c). Interestingly, when the direction of the magnetic field is reversed, the weakening behavior of the "X"-shaped constant energy contour also reverses between its two branches (yellow arrow in Fig. 2d).
To better illustrate this field-induced anisotropic spectral intensity change, spectral images along cuts across both K and K’ are extracted for a side-by-side comparison (Fig. 2e-h and Extended Data Fig. 1-2). The cuts are taken perpendicular to the –M mirror plane, where a mirror-symmetric spectral weight distribution is expected when there is no magnetic field. We note that the experimental geometry is chosen to preserve this mirror, as neither our beam polarization nor incident direction breaks this mirror symmetry (Fig. 2). We focus on the band as its spectral weight is discernible over a large energy range. For cut 1 within the first BZ, the ARPES spectra of the band on the left side () and those of the band on the right side () of the = 0 mirror plane are equivalent in the absence of a magnetic field (Fig. 2e). However, when an out-of-plane magnetic field (1.6 mT) is applied, the left branch of the band intensity becomes "weak" and the spectral weight becomes broader, as observed from the momentum distribution curve (MDC) in Fig. 2f. The spectral broadening is reversed with respect to the horizontal mirror plane when the magnetic field direction is reversed (Fig. 2g). This can be better seen in the magneto-dichroic ARPES signal shown in Fig. 2h, which we define as the difference in spectral intensity between the two spectra obtained in magnetic fields with opposite directions and the same strength. This magneto-dichroic signal can be well simulated by a simple selective momentum broadening of ARPES spectra based on the DFT calculated band structures along this cut (Extended Data Fig. 3). For cut 2, which lies outside the first BZ on the other side of the K–M–K BZ boundary, an opposite selective spectral weight broadening behavior is observed between and in a magnetic field compared with what is observed along cut 1 (Extended Data Fig. 1). The spectral evolution of the images along both cut 1 and cut 2 in an external magnetic field together complements the spectral response of the "X"-shaped energy iso-surface to a magnetic field, which directly breaks the –M mirror symmetry of the electronic states. We emphasize that this observation is an intrinsic electronic response of CsV3Sb5 to the external magnetic field, as the extrinsic momentum broadening effect is negligible at this small field and should be neither momentum selective nor dependent on the sign of the field Huang et al. (2023).
To pinpoint the precise symmetries that are broken by this anisotropic effect, we note that our measurement region also covers part of the second BZ, which can be mapped back to the first BZ via the crystalline translational symmetry. We also directly verified this by rotating the sample and mapping the different K/K’ pairs response to the magnetic field on the same sample (Extended Data Fig. 4). This leads to the illustration in Fig. 2i, where the broadened branches of the Fermi surface are depicted by the thicker and lighter-colored lines while the shaper branches are represented by the thinner and darker-colored lines. Comparing the behavior of the branches across the first BZ, we conclude that the inversion symmetry is preserved, whereas the sixfold (C6) rotational symmetry is broken. Moreover, since the effect changes sign when the field is reversed, the RSB is odd in field, which is indicative of piezomagnetism, i.e. a shear strain that is linear with the field Xing et al. (2024).
To further explore this effect, we performed temperature-dependent measurements of the spectral images of CsV3Sb5 at a fixed magnetic field (Fig. 2j-n and Supplementary Materials). Although the cut position (cut 3) slightly deviates from cut 1 in Fig. 2b, it still captures both the and bands (Fig. 2j). As the temperature increases, the originally asymmetric and bands in the CDW state restore their symmetry at higher temperatures. This is evident from the corresponding MDCs shown for eV: while the two peaks have distinct widths at K, they gradually restore to similar widths at higher temperatures (Fig. 2k). This evolution can also be visualized from the corresponding false-color map generated from the stacking of the temperature-dependent MDCs (Fig. 2l). We can further quantify this temperature dependence by fitting both the peak height and peak width of and from the MDCs. As seen from the resulting plots (Fig. 2m-n), both the peak intensity and width become asymmetric below , indicating that the symmetry breaking associated with the vanadium VHS bands is intrinsically linked to the CDW order.
Having demonstrated the magnetic-field-induced C6 symmetry-breaking of the electronic spectrum in the CDW state of CsV3Sb5, we further examine its detailed spectral weight evolution as a function of magnetic field (Fig. 3 and Supplementary Materials). The spectral images along the –K’ direction measured at different magnetic fields are obtained as a cut from the Fermi surface mappings using the DA30 deflector mode after correcting for the Lorentz rotation (Fig. 3a). In this measurement geometry, the aforementioned band is suppressed whereas the band acquires a more clear resolution. As expected from the analysis of the band, the spectral weight of the band significantly broadens at a positive magnetic field (2.1 mT), while the effect on the other bands are much less noticeable. To quantify this evolution, we extracted the MDCs at -0.2 eV from the –K’ spectral images (Fig. 3b). Three distinct peaks (P1, P2, and P3) are resolved: P1 originates from the electron band around the BZ center, P3 from the shallow Dirac point at K’, and P2 from the band. As the out-of-plane magnetic field is varied from negative to positive, peak P2 ( band) broadens significantly in comparison to the linewidths of peak P1 and peak P3 (Fig. 3d and top row of Fig. 3a). The evolution on the peak area follows a similar trend, though varying less in comparison to the peak width.
The odd-in-field response of the electronic structure to the out-of-plane field suggests signs of TRS-breaking. To understand the origin of the observed asymmetric spectral shape under a magnetic field, we carried out one-step model ARPES calculations using the spin-polarized relativistic Korringa-Kohn-Rostoker (SPRKKR) package to examine the effect of an external magnetic field on the photoemission spectral function on CsV3Sb5 Ebert et al. (2011) (see Methods and Extended Data Fig. 5). By calculating the spectral function of the non-CDW phase of CsV3Sb5 under an external field, we can simulate the photoemission spectra under a finite field for the simplest phase with broken TRS via the Zeeman effect without invoking a specific microscopic TRS-broken order. This set of calculations incorporates the magnetic field effects in all three parts of the photoemission matrix element—a Zeeman term for the initial state, the relativistic dipole operator, and the time-reversed spin-polarized low-energy electron diffraction final states, and is hence the simplest yet most unbiased way to examine the expectations from a generic TRS-broken state. As shown in Extended Data Fig. 5, the spectra near the K/K’ points exhibit asymmetric spectral shape both in peak height and widths when photoemitting under a large magnetic field. This is due to the combination of the field-induced band shift and spin-dependent multiple scattering in the photoemission process under a particular field direction. While at first sight this may seem consistent with our observations, the magnitude of the field needed for the Zeeman effect to be observable is five orders of magnitude larger than the field used in the magneto-ARPES measurements. Hence the electronic response experimentally observed here cannot arise from a Zeeman-like field-induced effect on the photoemission matrix elements, but must originate from an intrinsic TRS-breaking order parameter in CsV3Sb5.
III Field response of the Sb 5 band at
Next, we turn our attention to the electron pocket around the BZ center (Fig. 4), which arises from the Sb orbitals. When measured at zero field, as consistent with all previous reports Liu et al. (2021); Luo et al. (2022); Kang et al. (2022); Hu et al. (2022), the central electron pocket is nearly isotropic and circular inside the CDW state (Fig. 4a). This can be demonstrated by comparing the MDC along the vertical -K and horizontal -M directions, which shows that the separations of the two peaks are identical along the two orthogonal directions (Fig. 4b). However, when an out-of-plane magnetic field of -1.6 mT is applied, the circular Fermi pocket becomes elliptical (Fig. 4e), as can also be seen in the same comparison of the vertical (black) and horizontal (red) MDCs (Fig. 4f). Here we note that we can exclude the possibility that this Fermi surface distortion is due to potential misalignment between the sample and the magnetic coil, ellipticity of the beam, or spatial inhomogeneity of the field, as demonstrated via our extensive simulations (Supplementary Materials). Rather, the elliptical distortion of the central Fermi pocket under the application of a magnetic field suggests a field-induced RSB in CsV3Sb5 corresponding to broken symmetry. Beyond the Fermi surface, we also show the measured dispersions along the vertical and horizontal directions for both zero field and -1.6 mT (Fig. 4c,g). Their fitting shows that this elliptical distortion is not limited to the Fermi level but persists to a larger energy range of at least 250 meV below the Fermi level (Fig. 4d,h).
Another way to assess this anisotropy is to extract the fitted Fermi surface diameter as a function of the azimuthal angle (). As shown in Fig. 4i, the Fermi surface ellipticity for -1.6 mT is prominent compared to that at zero field at various temperatures. This trend is better seen by plotting the angle-dependent Fermi surface diameter in a magnetic field normalized by its zero-field value (Fig. 4j). We note that this angle-dependent normalized Fermi surface diameter can be well fitted by a phenomenological function of the form that captures the two-fold rotational symmetry, with denoting the Fermi surface anisotropy ratio and the angle denoting the deviation of the long axis of the elliptical Fermi surface from the BZ high symmetry –M (–K) direction. Intriguingly, we find that in the CDW phase in a magnetic field, the long axis of the elliptical Fermi surface deviates from the high-symmetry –K direction of the BZ (Fig. 4j and k). At the same time, the directional deviation of the elliptical Fermi surface also changes with temperature, going through zero at and reversing sign above (Fig. 4k). Additionally, the amplitude of the ellipticity of the Fermi surface remains relatively constant below and increases with increasing temperature above (Fig. 4l). We plot the temperature dependence of the pocket represented by the Fermi surface distortion together with the VHS spectral width difference extracted previously (Fig. 4l). Their behaviors are distinct across –the V 3-VHS spectral width difference onsets below while the electron pocket ellipticity is finite and largely constant below , and grows above up to the measured temperature of 125 K (Fig. 4m).
IV Discussion
Our magneto-ARPES measurements reveal distinct field responses, inside the CDW phase, of the V 3 orbitals that make up the bands associated with the VHS of the kagome lattice and the Sb orbitals that make up the band at the BZ center. The odd-in-field response of the VHS-bands breaks C6 rotational symmetry, thus implying that the CDW phase of CsV3Sb5 is piezomagnetic, i.e., that an out-of-plane magnetic field induces an in-plane distortion. Piezomagnetism, manifesting the intrinsic coupling between TRSB and RSB, was also inferred in a recent STM study on the related compound RbV3Sb5 Xing et al. (2024), which attributed this response to a so-called congruent flux phase with coexisting CDW and loop-current order. In this state, rotational symmetry is weakly broken already in the absence of a magnetic field, and the application of a magnetic field further breaks the remaining vertical mirror, leaving intact only the horizontal mirror. Our ARPES data at zero field, however, is symmetric, and the current magneto-ARPES data cannot resolve whether a vertical mirror plane is retained or not in the presence of a field. As discussed in the Methods section and in Extended Data Fig. 6-7, this could be due to the weak character of the zero-field symmetry-breaking or more likely domain averaging within the beamspot of the ARPES measurement. As our samples are glued down, they are not free from residual strain. Considering that our beamspot is much larger than the congruent flux order domains, in the absence of field, the time-reversed pairs of loop current domains would both appear, averaging out in our ARPES signal. With the application of a field and also in the presence of residual strain, one type of time-reversed loop current domain would dominate, revealing the observed electronic response. To distinguish these scenarios from the one where symmetry is preserved at zero-field, future nano-focused synchrotron-based magneto-ARPES measurements will be helpful. Regardless of its microscopic origin, the piezomagnetic response, which implies that the CDW state spontaneously breaks TRS, is robust and strongly manifested in the V bands that give rise to the VHS, which are often associated with the emergence of unconventional correlated states in the kagome lattice Kiesel et al. (2013). A field-induced breaking of symmetry is also manifested in the bands made up of the Sb orbitals, which have been proposed to be essential for superconductivity Ritz et al. (2023). Our measurements on this Fermi pocket ellipticity around the BZ center, however, do not exhibit a clear sign change upon reversing the magnetic field (Extended Data Fig. 8), in contrast to the clear odd-in-field behavior displayed by the anisotropy of the V bands.
It is interesting to note that the field-response of the V 3 bands vanishes at while that of the Sb bands persists above . The origin of the latter behavior and its connection to the CDW order will require further studies. Interestingly, a recent magneto-torque measurement reported rotational symmetry breaking above that was interpreted to also break inversion symmetry Asaba et al. (2024), which is not the case in our magneto-ARPES data. One possibility is that strong fluctuations persist above and are thus “picked up” by the magnetic field, similarly to how small residual strain induces robust nematic order in the iron-based superconductors due to an enlarged nematic susceptibility Fernandes et al. (2014). Indeed, it has been demonstrated that CsV3Sb5 is extremely sensitive to both magnetic field and strain Guo et al. (2024a), and our glued sample is not free from random residual strain, which may combine with the external magnetic field to induce an ordered state. Alternatively, the magnetic field could also enhance the separation between the onsets of the CDW and loop-current order parameters in certain loop-current configurations Christensen et al. (2022); Wagner et al. (2023); Tazai et al. (2024). In either scenario, however, it is puzzling why this behavior is seen only in the Sb bands and why the anisotropy is larger outside the CDW phase than inside the CDW phase. It remains to be seen whether future experiments can map the electronic response of the distinct momentum-space regions to decoupled strain and magnetic fields, which will be helpful in disentangling different microscopic mechanisms. The new capability of magneto-ARPES demonstrated here, which enables the mapping of the spectral response in a momentum-resolved way in the presence of a tunable finite magnetic field, opens up that possibility.